A boy walks at a speed of 4km/h. How much time does he take to walk a distance of 20km?
1
Expert's answer
2018-06-13T08:31:08-0400
We can find how much time the boy need to walk the distance of 20 km from the formula:
t=s/v=(20 km)/(4 km/h)=5 h.